OpenStudy (anonymous):

Solve the equation for V^2

OpenStudy (anonymous):

\[\left(\begin{matrix}P^1V^1 \\ T^1\end{matrix}\right)=\left(\begin{matrix}P^2V^2 \\ T^2\end{matrix}\right)\]